# pre-trained model AST

### v1

@@ 4. System and Threat Model
Our motivational study shows the need for a new defense that is both usable and secure. Here, we model the system we aim to protect, and the adversary to be thwarted:
@@
We focus on the threat of pickle program code execution, and exclude other threats from untrusted ML models. Other ML supply chain attacks are orthogonal and can be approached with layered defenses, like manipulating model weights to insert “backdoors” (Clifford et al., 2024) and compromising other model components. We don’t consider other forms of attacks that manipulate pickle programs, e.g., for denial of service (coldwaterq, 2024) due to their weaker attack primitives.
@@ 5. PickleBall Design and Implementation
[…]
@@
[…]
@@
PickleBall is designed to protect applications that use libraries to load untrusted pickle models. The desired system guarantee is that PickleBall raises a security exception when an adversary invokes an unnecessary callable during model loading, while transparently loading benign models. The idea behind PickleBall is to first generate a policy describing a minimal set of operations (i.e., the callables that need to be imported and invoked) for instantiating a given model object during an offline policy generation phase (§5.1), and then to enforce the generated policy during unpickling, rejecting spurious operations performed by malicious models (§5.2). Figure˜5 provides an overview of the aforementioned PickleBall components. PickleBall guarantees that, given correct AST and type information, PickleBall raises a security exception when the adversary invokes an unnecessary callable (§5.3). We implement the design of PickleBall as a software artifact (§5.4).
@@ 5.1. Policy Generation
PickleBall’s policy generation component is designed to automatically create a policy that describes the set of operations (i.e., the set of callables imported and invoked) permitted when loading a model for a given loading API. The policy is generated before loading the untrusted pickle model, and restricts the loading behaviors to only those that are necessary for the library API.
@@
The challenging task of policy generation is determining which callables are needed to instantiate an object. Python is a dynamic language that permits a single variable to receive different types at various program paths, and for objects to receive new attribute variables after initialization. Identifying all attribute types requires a path-sensitive analysis of the object creation code up to the point that the object is serialized. Further, in the ML setting, the code that creates the object (model) is not provided: the model saving program is not provided alongside the serialized model, since it s used to trained the model; the goal of sharing the serialized model to obviate the need for executing a training program.
However, we recognize that we can generate approximate policies for serialized models by analyzing the class definition of the model, even without access to the model saving program. The majority of the class instance attribute information is contained in the class definition, in the shared library code, rather than in the model saving application. Intuitively, for the object to be reusable, it has to conform to an expected interface that is defined in the library.
@@
    Otherwise: add the class to the allowed imports sets; add all sub-classes of the class to the candidate queue; add all types of the class’s attributes (including attributes inherited) to the candidate queue.
@@
PickleBall’s static analysis cannot produce perfectly sound and precise policies due to fundamental challenges in statically analyzing Python code, which is dynamically typed (Xu et al., 2016; Aycock, 2000; Gorbovitski et al., 2010). When recovered type information is over-approximate, PickleBall produces policies that contain more callables than is strictly necessary. Python’s dynamic features, like dynamic typing and runtime attribute manipulation, can lead to unaccounted data dependencies during type recovery, resulting in missing type information and therefore in policies that incorrectly exclude callables. To account for these potential errors, we design for security by separately enforcing allowed import and allowed invocations, and for robustness with a lazy enforcement mechanism (§5.2). We discuss how static analysis limitations affect the whole-system analysis in Section˜5.3.
@@ 5.2. Policy Enforcement
PickleBall’s policy enforcement module is designed to protect the model loading application during loading. It is a drop-in replacement for the system pickle module. It restricts the behavior of the Pickle Machine operations that import, initialize, and invoke Python callables, so that only the callables permitted by the model loading policy are accessible to the pickle program in the model.
@@ 5.3. Security Guarantees and Limitations
Whole-system Guarantees: PickleBall’s design thwarts the adversary described in our Threat Model (§4) by restricting the Pickle Machine’s access to Python callables. The design guarantees that when provided a correct AST with type information for a Python class definition and its dependencies, PickleBall (1) raises a security exception when the adversary invokes a spurious callable, while (2) successfully instantiating any object that respects the attribute and type information defined class definition. In our evaluation of PickleBall, we show that existing state-of-the-art static analysis tools provide sufficient AST and type-recovery information for practical use (§6).
Policy Generation Guarantees: When the PickleBall is provided a correct AST with type information, it is guaranteed to output a loading policy that includes all allowed imports and allowed invocations that can appear in when an object is saved in the pickle format, provided that the object is not manipulated to add attribute types outside of its object prototype.
@@ 5.4. Implementation
PickleBall is implemented in a total of $\sim$}1,300 Scala lines of code (LoC) and $\sim$}300 Python LoC divided between two primary components: a static program analysis that builds upon the Joern framework (Yamaguchi et al., 2014); and a dynamic loader that modifies the existing Pickle Machine. In the static analysis, $\sim$}700 Scala LoC implement Algorithm˜1, $\sim$}600 Scala LoC extend and fix Joern features, and $\sim$}200 Python LoC integrate components. In the loader, $\sim$}100 Python LoC modify the Pickle Machine to implement lazy policy enforcement. Joern provides a program analysis platform for PickleBall by generating a Code Property Graph (CPG) with recovered type information for the target code; we extend Joern to improve type recovery features and class inheritance tracking (for more details, see Appendix˜C). Our analysis then queries Joern’s AST nodes for the relevant information.
Limitations: PickleBall inherits some limitations from the Joern program analysis framework. The limitations include that PickleBall:
@@ 6.1. Constructing an Evaluation Dataset
Benign Models and Trusted Libraries: Our dataset must represent the typical model reuse tasks of loading and inference; therefore, we need a set of benign models and the libraries used to load and interact with them. We first searched for libraries that meet three criteria: (1) they load pickle models; (2) they have a model class type that is returned by a loading API; and (3) if a foundational library (e.g., PyTorch or transformers) type is used, the custom class adds new attributes to the type. These criteria are motivated by PickleBall’s purpose: to restrict the pickle operations permitted by a library loading API based on analysis of the intended class type.
@@
We identified candidate libraries by identifying popular pickle models on Hugging Face and working backwards. We first searched Hugging Face programmatically for pickle models, ordered by monthly download rate, that had model loading documentation directing users to a model loading library API. We manually reviewed the top 400 models (approximately 2 hours of review time) to determine whether the identified libraries meet our criteria; this resulted in 16 accepted libraries. All libraries and their associated version information are listed in Appendix˜D (Table˜3).
@@
Then, we identified candidate models associated with each library. We again queried Hugging Face to identify models associated with the library, either directly (as a piece of repository metadata) or by mention in the model documentation or name. We collect models with $\geq$ 100 monthly downloads at time of collection. In total, we accumulated 252 models produced by 16 different libraries. All collected models are listed in Appendix˜D (LABEL:TABLE:model-versions).
@@
We acknowledge that these models could themselves be malicious. We partially mitigate this by sampling from the most frequently downloaded models and libraries, checking model scanner indications, and investigating unexpected callables when identified by the restricted model loaders by checking for their origin in library source code.
@@
Malicious Models: Our dataset must also represent the models created by our intended adversary (cf. our threat model — §4); therefore, we need a set of malicious pickle models. We first collected 82 malicious models and pickle programs that were identified on Hugging Face by two state-of-the-art model scanners (Casey et al., 2024; Zhao et al., 2024). We add our 2 malicious models constructed to bypass scanners (see Appendix˜B), for a total of 84. All malicious models contain pickle programs with payloads that import and invoke Python spurious functions; payload behaviors include accessing sensitive files, making network connections, and creating reverse shells.
@@
We acknowledge that these models do not represent the complete set of malicious model behaviors; it is a best-effort collection of real-world pickle model malware that represents today’s attackers. We aim to thwart the adversary that executes arbitrary Python functions during pickle loading, and the existing attacks all exercise this feature.
@@
Test Harnesses: PickleBall protects loading applications from untrusted models during and after loading; therefore, we need a representative set of loading programs to secure. We create one test harness for each library in our dataset; the harness loads a model using the library API and performs an inference task.
@@
Library Pre-processing: We pre-process the libraries before analyzing them to account for limitations in the static analysis framework (§5.4) and improve the correctness of the AST. We make manual source code modifications (<<10 LoC) when the library class uses newer Python features of Python that Joern’s front-end parser does not support, like generic type inheritance and type variables, or when the analysis misidentifies an imported library alias. For libraries that provide type hints in docstrings, which Joern does not parse, we copy (but do not modify) the hints as type annotations (<<100 LoC). We manually copy dependencies into the analysis scope when discovered during policy generation. Because model loading policies are compositional, we pre-compute policies for some frequently reused dependencies, including classes from the Python standard library and PyTorch, and save them in a class “cache” for PickleBall to access when it recognizes one of the classes in its analysis. Due to the complexity and prevalence of dynamically dispatched and C code implementations in PyTorch, we supplement our analysis of PyTorch modules with the weights-only unpickler in the class cache.
@@ 6.2. RQ1: Malicious Model Blocking
PickleBall must protect loading programs from pickle models with malicious payloads. To evaluate this, we assess whether any of our test harness PickleBall policies permit any malicious model executions.
@@ 6.2.1. Methods
For each harness program in our dataset, PickleBall generates a loading policy by analyzing the library and associated model class. We use PickleBall to enforce the generated policy while the harness attempts to load all malicious models in our dataset. For each malicious model, we consider the model blocked if PickleBall raises a security exception during model loading or inference, preventing the payload from executing.
@@ 6.3. RQ2: Benign Model Loading
PickleBall must enable users to load and perform tasks with benign models. Its policies must not be so restrictive that the models are unusable. We therefore evaluate PickleBall’s policies for loading and correctly using the benign models in our dataset (§6.3.1 and §6.3.2). To ensure robustness of the loaded model despite lazy enforcement, we perform a dataset test of successfully loaded models with stub objects (§6.3.3 and §6.3.4).
@@ 6.3.2. Results
PickleBall generates policies that, when enforced, correctly load and execute 79.8% of benign models in the dataset (Table˜1). In most cases, the policies contain all callables (Table˜1 – Imports and Invocations Allowed) that are seen in the model traces (Imports and Invocations Observed). In some cases, PickleBall’s policies do not include callables that are included in the models (flair, PyAnnote, YOLOv5, and YOLOv11), resulting in the creation of a stub object, and which is occasionally invoked, resulting in a security violation.
We investigated the failed models to determine their causes:
@@
  • Follow-on pickle loading: PickleBall fails to load two models from the MeloTTS library after they have been loaded, due to additional pickle loading during inference. PickleBall’s policy includes all callables needed to load the MeloTTS models. However, during inference, an additional pickle model is loaded; this model invokes a disallowed callable, resulting in a security violation.
@@ 6.3.3. Lazy Enforcement Robustness – Methods
Due to PickleBall’s lazy enforcement, models can load successfully despite being instantiated with stub objects. To further ensure that our benign models are robustly instantiated for inference, we evaluate these models with a more rigorous test suite of inputs during inference. We investigated the libraries that successfully load models with stub objects, i.e., Flair, PyAnnote, YOLOv5, and YOLOv11.
For each library, we find an extensive evaluation dataset to test each loaded model with: for Flair, we used various Named Entity Recognition and Universal Dependencies (Dependencies, 2025) datasets that come pre-packaged with the Flair library (Flair, 2025); for PyAnnote, we used the AISHELL-4 speech dataset (Fu et al., 2021); for YOLOv5 and YOLOv11, we used the 2017 Test Images Common Objects in Context dataset (Lin et al., 2014).
@@ 6.4. RQ3: Performance
PickleBall must be fast enough for practical use in developer and user tasks. We analyze two aspects of PickleBall’s performance: (1) the time to generate a policy for a class, which is an offline, one-time analysis cost (see §5.1), and (2) the additional runtime overhead of enforcing a policy to load and use a model, compared to the regular Pickle Machine.
@@ 6.4.1. Methods
To measure the time to generate policies, we execute PickleBall’s policy generator three times for each library in our dataset and compute the average between the three. We measure the real time using the Python time library. We run this experiment on a laptop with a 14-core Intel i7 CPU and 32GB of RAM (Linux Mint 22), representing a commodity developer environment.
To measure the additional runtime overhead of PickleBall’s policy enforcer, we isolate and record the time each harness program spends invoking the pickle load function during model loading. We first execute harness program with the unrestricted Pickle Machine environment to load a benign model. Then, we perform the same execution with PickleBall enabled. For fair comparison, we ensure that the unrestricted environment always uses the Python implementation of the Pickle Machine, instead of an optimized C implementation. We run this experiment on a server with a 32-core AMD EPYC 7502 processor and 256GB of RAM (Ubuntu 24.04); this is used for the attached hard-drive space for interacting with the hundreds of models in our dataset.

### v2

@@ 4. System and Threat Model
Our motivational study shows the need for a new defense that is both usable and secure. Here, we model the system we aim to protect and the adversary to thwart:
@@
We focus on the threat of pickle program code execution, and exclude other threats from untrusted ML models. Other ML supply chain attacks, like manipulating model weights to insert “backdoors” (Gu et al., 2019; Clifford et al., 2024), are orthogonal and can be approached with layered defenses. We do not consider other forms of attacks that manipulate pickle programs, e.g., for denial of service (coldwaterq, 2024) due to their weaker attack primitives.
@@ 5. PickleBall Design and Implementation
Policy Generationlib.pyLibrary Code Model Class Identifier Policy Generator Static Analysis Model Loading Policy Allowed Imports Allowed Callables app.py Model LoadingPickleBall LoaderPickle Program GLOBAL ModelClass
@@
... REDUCE foo(arg) data.pkl Allowed Import? Allowed Callable? ModelClass Stub Object foo(arg) model = library.load(’data.pkl’) model.infer() . Exception ModelClass.infer()
$p$
@@
PickleBall is designed to protect applications that use libraries to load untrusted pickle models. The desired system guarantee is that PickleBall raises a security exception when an adversary invokes an unnecessary callable during model loading, while transparently loading benign models. The idea behind PickleBall is to first generate a policy describing a minimal set of operations (i.e., the callables that need to be imported and invoked) for instantiating a given model object (§5.1), and then to enforce the generated policy during unpickling, rejecting spurious operations performed by malicious models (§5.2). Figure˜5 provides an overview of the aforementioned PickleBall components. PickleBall guarantees that, given correct AST and type information, PickleBall raises a security exception when the adversary invokes an unnecessary callable (§5.3). We implement the design of PickleBall as a software artifact (§5.4).
@@ 5.1. Policy Generation
PickleBall’s policy generation component is designed to automatically create a policy that describes the set of operations (i.e., the imported and invoked callables) permitted when loading a model of a particular library class. The policy is generated before loading the untrusted pickle model, and restricts the loading behaviors to only those that are necessary for the library API.
@@
The challenging task of policy generation is determining which callables are needed to instantiate an object. Python is a dynamic language that permits a single variable to receive different types at various program paths, and for objects to receive new attribute variables after initialization. Identifying all attribute types requires a path-sensitive analysis of the object creation code up to the point that the object is serialized. Further, in the ML setting, the code that creates the model object is not provided: the model saving program that trained the model is not published along with the model.
However, we recognize that we can generate approximate policies for serialized models by analyzing the class definition of the model in the library source code, even without access to the model saving program. The majority of the class instance attribute information is contained in the class definition, rather than in the model saving application. Intuitively, for the object to be reusable, it has to conform to an expected interface that is defined in the library.
@@
    Otherwise: add the class to the allowed imports set; add all sub-classes of the class to the candidate queue; add all types of the class’s attributes (including attributes inherited) to the candidate queue.
@@
PickleBall’s static analysis cannot produce perfectly sound and precise policies due to fundamental challenges in statically analyzing Python code, which is dynamically typed (Xu et al., 2016; Aycock, 2000; Gorbovitski et al., 2010). When recovered type information is over-approximate, PickleBall produces policies that contain more callables than is strictly necessary. Under-approximate type information may result from Python’s dynamic features, like dynamic typing and runtime attribute manipulation, producing unaccounted data dependencies during type recovery. The missing type information creates policies that incorrectly exclude callables. To account for these potential errors, we design for security by separately enforcing allowed import and allowed invocations, and for robustness with a lazy enforcement mechanism (§5.2). We discuss how static analysis limitations affect the whole-system analysis in Section˜5.3.
@@ 5.2. Policy Enforcement
PickleBall’s policy enforcement module is designed to protect the model loading application during loading. It is a drop-in replacement for the system pickle module. It restricts the behavior of the PM operations that import, initialize, and invoke Python callables, so that only the callables permitted by the model loading policy are accessible to the pickle program in the model.
[…]
@@ 5.3. Security Guarantees and Limitations
Whole-system Guarantees: PickleBall’s design thwarts the adversary described in our Threat Model (§4) by restricting the PM’s access to Python callables. The design guarantees that when provided a correct AST with type information for a Python class definition, PickleBall (1) raises a security exception when the adversary invokes a spurious callable, while (2) successfully instantiating any object that respects the attribute and type information defined class definition. In our evaluation of PickleBall, we show that existing state-of-the-art static analysis tools provide sufficient AST and type-recovery information for practical use (§6).
Policy Generation Guarantees: When PickleBall is provided a correct AST with type information, it is guaranteed to output a loading policy that includes all allowed imports and allowed invocations that can appear in when an object is saved in the pickle format, provided that the object is not manipulated to add attribute types outside of its object prototype.
@@ 5.4. Implementation
PickleBall is implemented in a total of $\sim$}1,300 Scala lines of code (LoC) and $\sim$}300 Python LoC divided between two primary components: a static program analysis that builds upon the Joern framework (Yamaguchi et al., 2014); and a dynamic loader that modifies the existing Pickle Machine. In the static analysis, $\sim$}700 Scala LoC implement Algorithm˜1, $\sim$}600 Scala LoC extend and fix Joern features, and $\sim$}200 Python LoC integrate components. In the loader, $\sim$}100 Python LoC modify the Pickle Machine to implement lazy policy enforcement. Joern provides a program analysis platform for PickleBall by generating a Code Property Graph (CPG) with recovered type information for the target code; we extend Joern to improve type recovery features and class inheritance tracking (for more details, see Appendix˜C). Our analysis queries Joern’s AST nodes for the relevant information.
Limitations: We inherit some limitations from the Joern program analysis framework. The limitations include that PickleBall:
@@ 6.1. Constructing an Evaluation Dataset
Benign Models and Trusted Libraries: Our dataset must represent how users typically load and use models; therefore, we need a set of benign models and the libraries used to load and interact with them. We first searched for libraries that meet three criteria: (1) they load pickle models; (2) they have a model class type that is returned by a loading API; and (3) if a foundational library (e.g., PyTorch or transformers) type is used, the custom class adds new attributes to the type. These criteria are motivated by PickleBall’s purpose: to restrict the pickle operations permitted by a library loading API based on analysis of the intended class type.
@@
We identified candidate libraries by identifying popular pickle models on Hugging Face and working backwards. We first searched Hugging Face programmatically for pickle models, ordered by monthly download rate, that had model loading documentation directing users to a model loading library API. We manually reviewed the top 400 models (approximately 2 hours of review time) to determine whether the identified libraries meet our criteria; this resulted in 16 accepted libraries. All libraries and their associated version information are listed in Table˜3, Appendix˜D.
@@
Then, we identified candidate models associated with each library. We again queried Hugging Face to identify models associated with the library, either directly (as a piece of repository metadata) or by mention in the model documentation or name. We collected models with $\geq$ 100 monthly downloads at time of collection. In total, we accumulated 252 models produced by 16 different libraries. All collected models are listed in LABEL:TABLE:model-versions, Appendix˜D.
@@
We acknowledge that these models could themselves be malicious. We partially mitigate this by sampling from the most frequently downloaded models and libraries, checking model scanner indications, and manually investigating unexpected callables when restricted model loaders identify them.
@@
Malicious Models: Our dataset must also represent the models created by our intended adversary (cf. our threat model — §4); therefore, we need a set of malicious pickle models. We first collected 82 malicious models and pickle programs that were identified on Hugging Face by two state-of-the-art model scanners (Casey et al., 2024; Zhao et al., 2024). We add our 2 malicious models constructed to bypass scanners (see Appendix˜B), for a total of 84. All malicious models contain pickle programs with payloads that import and invoke Python spurious functions; payload behaviors include accessing sensitive files, making network connections, creating reverse shells, among others.
@@
We acknowledge that these models do not represent the complete set of malicious behaviors; it is a best-effort collection of real-world pickle model malware to represent today’s attackers. We aim to prevent adversaries that execute arbitrary Python functions during pickle loading, and the collected models all exercise this feature.
@@
Test Harnesses: PickleBall protects programs that load untrusted models; therefore, we need a representative set of loading programs to secure. We create one test harness for each library in our dataset; the harness loads a model using the library API and performs an inference task.
@@
Library Pre-processing: We pre-process the libraries before analyzing them to account for limitations in the static analysis framework (§5.4) and improve the correctness of the AST. We make manual source code modifications (<<10 LoC) when the library class uses newer Python features of Python that Joern’s front-end parser does not support, like generic type inheritance and type variables, or when the analysis misidentifies an imported library alias.
For libraries that provide type hints in docstrings, which Joern does not parse, we copy (but do not modify) the hints as type annotations (<<100 LoC). We manually copy dependencies into the analysis scope when discovered during policy generation. Because model loading policies are compositional, we pre-compute policies for some frequently reused dependencies, including classes from the Python standard library and PyTorch, and save them in a class “cache” for PickleBall to access when it recognizes one of the classes in its analysis. Due to the complexity and prevalence of dynamically dispatched and C code implementations in PyTorch, we supplement our analysis of PyTorch modules with the weights-only unpickler policy in the class cache.
@@ 6.2. RQ1: Malicious Model Blocking
PickleBall must protect loading programs from pickle models with malicious payloads. To evaluate this, we assess whether any of the generated PickleBall policies allow malicious model executions.
@@ 6.2.1. Methods
For every harness program in our dataset, PickleBall generates a loading policy by analyzing the library and associated model class. We use PickleBall to enforce the generated policy while the harness attempts to load all malicious models in our dataset. For each malicious model, we consider the model blocked if PickleBall raises a security exception during model loading or inference, preventing the payload from executing.
@@ 6.3. RQ2: Benign Model Loading
PickleBall must let users load and perform tasks with benign models. Its policies must not be so restrictive that the models are unusable. We therefore evaluate PickleBall’s policies for loading and correctly using the benign models in our dataset (§6.3.1 and §6.3.2). To ensure robustness of the loaded model despite lazy enforcement, we further test the successfully loaded models that contain stub objects (§6.3.3 and §6.3.4).
@@ 6.3.2. Results
PickleBall generates policies that, when enforced, correctly load and execute 79.8% of benign models in the dataset (Table˜1). In most cases, the policies contain all callables (Table˜1 – Imports and Invocations Allowed) that are seen in the model traces (Imports and Invocations Observed). In some cases, PickleBall’s policies do not include callables that are included in the models (flair, PyAnnote, YOLOv5, and YOLOv11), resulting in the creation of a stub object that is occasionally invoked, hence leading to a security violation. We investigated the failed models/cases to determine their causes:
@@
  • Follow-on pickle loading: PickleBall fails to load two models from the MeloTTS library after they have been loaded, due to additional pickle loading during inference. PickleBall’s policy includes all callables needed to load the MeloTTS models. However, during inference, an additional pickle model is loaded that invokes a disallowed callable, resulting in a security violation.
@@ 6.3.3. Lazy Enforcement Robustness – Methods
Due to PickleBall’s lazy enforcement, models can load successfully with incomplete attributes. To further ensure that the benign models are robustly instantiated with all attributes needed for inference, we evaluate these models with a more rigorous test suite of inputs. We investigated the libraries that successfully load models with stub objects, i.e., Flair, PyAnnote, YOLOv5, and YOLOv11.
For each library, we find an extensive evaluation dataset to test each loaded model: for Flair, we used various Named Entity Recognition and Universal Dependencies (Dependencies, 2025) datasets that come pre-packaged with the Flair library (Flair, 2025); for PyAnnote, we used the AISHELL-4 speech dataset (Fu et al., 2021); for YOLOv5 and YOLOv11, we used the 2017 Test Images Common Objects in Context dataset (Lin et al., 2014).
@@ 6.4. RQ3: Performance
PickleBall must be fast enough for practical use in developer and user tasks. We analyze two aspects of PickleBall’s performance: (1) the time to generate a policy for a class, which is an offline, one-time analysis cost (§5.1), and (2) the additional runtime overhead of enforcing a policy to load and use a model, compared to the regular Pickle Machine.
@@ 6.4.1. Methods
To measure the time to generate policies, we execute PickleBall’s policy generator three times for each library in our dataset and compute the average between the three. We measure the real time using the Python time library. We run this experiment on a laptop with a 14-core Intel i7 CPU and 32GB of RAM, representing a commodity developer environment.
[…]